Find the area of the shaded region. The graph to the right depicts IQ scores of​ adults, and those scores are normally distributed with a mean of 100 and a standard deviation of 15

The left boundary is at 80 and the right boundary is at 125
Answer #1

Solution :

Given that ,

mean = = 100

standard deviation = = 15

P(80 < x < 125) = P((80 - 100)/ 15) < (x - ) / < (125 - 100) / 15) )

= P(-1.33 < z < 1.67)

= P(z < 1.67) - P(z < -1.33)

= 0.9525 - 0.0918

= 0.8607

The left boundary is at 80 and the right boundary is at 125 = 0.8607